### v1

@@ Acknowledgements.
DM acknowledges useful conversations with Márton Balázs and funding from an EPSRC DTP studentship, grant number EP/T517872/1. LG, DM and FB thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programmes Mathematics of Movement: an interdisciplinary approach to mutual challenges in animal ecology and Cell Biology and Stochastic Systems for Anomalous Diffusion, where work was, respectively, initiated and concluded supported by EPSRC grants EP/R014604/1 and EP/Z000580/1. LG also acknowlegdes funding from NERC grant NE/W00545X/1. FB and PF also acknowledge the funding from the Ministry of Science, Innovation, and Universities in Spain through grant no. PID2021-122893NB-C21.

### v2

@@ Acknowledgements.
DM acknowledges useful conversations with Márton Balázs and funding from an EPSRC DTP studentship, grant number EP/T517872/1. LG, DM and FB thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programmes Mathematics of Movement: an interdisciplinary approach to mutual challenges in animal ecology and Cell Biology and Stochastic Systems for Anomalous Diffusion, where work was, respectively, initiated and concluded supported by EPSRC grants EP/R014604/1 and EP/Z000580/1. LG also acknowlegdes funding from NERC grant NE/W00545X/1 and support by a grant from the Simons Foundation. FB and PF also acknowledge the funding from the Ministry of Science, Innovation, and Universities in Spain through grant no. PID2021-122893NB-C21.
Appendix 1: Definition of Matrices in Original Master Equation
The matrices in Eq. (2) are defined as follows:
$\mathbb{A}$ =$=\{\mathdutchcal{a}_{i,j}\}\text{ where }\mathdutchcal{a}_{3,4}=l,\;\mathdutchcal{a}_{3,5}=r,\;\mathdutchcal{a}_{3,6}=b,$ (23)
$\mathbb{B}$ =$=\{\mathdutchcal{b}_{i,j}\}\text{ where }\mathdutchcal{b}_{6,1}=l,\;\mathdutchcal{b}_{6,2}=r,\;\mathdutchcal{b}_{6,3}=b,$
$\mathbb{C}$ =$=\{\mathdutchcal{c}_{i,j}\}\text{ where }\mathdutchcal{c}_{5,1}=r,\;\mathdutchcal{c}_{5,2}=b,\;\mathdutchcal{c}_{5,3}=l,$
$\mathbb{D}$ =$=\{\mathdutchcal{d}_{i,j}\}\text{ where }\mathdutchcal{d}_{2,4}=r,\;\mathdutchcal{d}_{2,5}=b,\;\mathdutchcal{d}_{2,6}=l,$
$\mathbb{E}$ =$=\{\mathdutchcal{e}_{i,j}\}+c\mathbb{I}\text{ where }\mathdutchcal{e}_{4,1}=\mathdutchcal{e}_{1,4}=b,\;\mathdutchcal{e}_{4,2}=\mathdutchcal{e}_{1,5}=l,\;\mathdutchcal{e}_{4,3}=\mathdutchcal{e}_{1,6}=r,$
where $\mathbb{I}$ is the $6\times 6$ identity matrix and where all other indices are zero.